OpenStudy (anonymous):

Solve the inequality. Graph the solution set. 2r – 9(Greater or equal to sign) – 6

OpenStudy (lena772):

2r-9>=-6 add 6 to each side 2r-3>=0 2r>=3 r>=2/3

OpenStudy (anonymous):

i REALLY have noo idea!

OpenStudy (lena772):

http://www.wolframalpha.com/input/?i=2r-9%3E%3D-6

OpenStudy (lena772):

r>=3/2 ***

OpenStudy (anonymous):

Sorry but that isnt a choice... the > is the one with line below it

OpenStudy (anonymous):

NOT answers... click the picture i attached... its the full question with the choices...

OpenStudy (lena772):

3/2 = 1.5 or 1 1/2

OpenStudy (lena772):

So it's between A & C... look at the URL i posted

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Thanks so much

OpenStudy (lena772):

Yes!

OpenStudy (anonymous):

Fanned =] thanks